OpenStudy (anonymous):

F(x,y) = x^3 - y^3 - 2xy + 6 the critical points i obtained are (0,0), (2/3,2/3) and (2/3, -2/3), but the solution for the critical points are (0,0) and (-2/3,2/3). did i made any mistake somewhr?

OpenStudy (anonymous):

*solution gvs (0,0) and (-2/3,2/3)

OpenStudy (anonymous):

mayb i should provide my workings also?

OpenStudy (experimentx):

Must be some calculation error http://www.wolframalpha.com/input/?i=3x%5E2+-+2y+%3D+0%3B+-3y%5E2+-+2x+%3D+0

OpenStudy (anonymous):

omg.. now i see it. LOL

OpenStudy (anonymous):

its the diff wrt y, missed out the - ._.

OpenStudy (anonymous):

sry for the fuss

OpenStudy (anonymous):

erm, so when\[x=-\frac{2}{3},\]\[y ^{2}=-\frac{2}{3}(-\frac{2}{3})\]why y=2/3 only?

OpenStudy (anonymous):

shouldnt it be + - ?

OpenStudy (anonymous):

@arnab09 whats your opinion on the y=2/3 only?